OpenStudy (jocelynstalnaker):

If f(x) = 2x^2 - 30, find f(4). 35 -14 2 17

OpenStudy (jocelynstalnaker):

@marissalovescats

OpenStudy (marissalovescats):

Plug in 4 for x

OpenStudy (jocelynstalnaker):

f(4)= 2*4^2-30

OpenStudy (marissalovescats):

Yes

OpenStudy (acxbox22):

f(4)=2(4)^2-30 f(4)=2(16)-30 f(4)=32-30 f(4)=???

OpenStudy (jocelynstalnaker):

I the answer 2?

OpenStudy (jocelynstalnaker):

Ye

OpenStudy (jocelynstalnaker):

Thak you!

OpenStudy (acxbox22):



OpenStudy (marissalovescats):

Yes it's 2. You're welcome :) Great job!

OpenStudy (jocelynstalnaker):

Lol, could y'all help me with one more?

OpenStudy (marissalovescats):

Yeah I can

OpenStudy (jocelynstalnaker):

I dont understand this problem The first five terms of a sequence are shown below: 4, 7, 10, 13, 16 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? f(n) = 3n - 1; n ≥ 1 f(n) = 3n + 1; n ≥ 1 f(n) = 4n - 1; n ≥ 1 f(n) = 4n + 1; n ≥ 1

OpenStudy (marissalovescats):

So this is an arithmetic sequence and there is a formula for it, it is an=a1+(n-1)d Where a1 is the first term and d is the difference

OpenStudy (jocelynstalnaker):

I dont know if its either B or C

OpenStudy (marissalovescats):

The first term in the sequence they gave you was 4 then 7,10,13,16,etc. So you can tell the difference is that you are adding 3

OpenStudy (marissalovescats):

So we plug it in